# Are Triggers Needed for Document-Level Event Extraction?

### v1

@@ Abstract
Most existing work on event extraction has focused on sentence-level texts and presumes the identification of a trigger-span — a word or phrase in the input that evokes the occurrence of an event of interest. Event arguments are then extracted with respect to the trigger. Indeed, triggers are treated as integral to, and trigger detection as an essential component of, event extraction. In this paper, we provide the first investigation of the role of triggers for the more difficult and much less studied task of document-level event extraction. We analyze their usefulness in multiple end-to-end and pipelined neural event extraction models for three document-level event extraction datasets, measuring performance using triggers of varying quality (human-annotated, LLM-generated, keyword-based, and random). Our research shows that trigger effectiveness varies based on the extraction task’s characteristics and data quality, with basic, automatically-generated triggers serving as a viable alternative to human-annotated ones. Furthermore, providing detailed event descriptions to the extraction model helps maintain robust performance even when trigger quality degrades. Perhaps surprisingly, we also find that the mere existence of trigger input, even random ones, is important for prompt-based LLM approaches to the task.
**footnotetext: These authors contributed equally to this work
@@ 1 Introduction
Figure 1: Document-level event extraction example from MUC. Two events should be extracted from the given text, an Attack (doc on left) and a Kidnapping (doc on right). Note that only one mention (among many coreferent ones) per event argument is shown. Some event arguments are empty and some might have multiple fillers. The (duplicated) document also depicts event triggers obtained from different annotation sources (labeled with pink background); more details in Section 3.
@@
In recent years, however, there has been increasing interest in document-level event extraction: given a document $d$, a set of event types $E$, and their associated arguments $A_{e\in E}$, produce one event output template for each relevant event described in $d$ with event arguments recognized at the entity level, i.e. via a single representative mention. [^1] Other names for the task include template filling Jurafsky and Martin (2008), generalized template extraction (Chen et al., 2023), N-ary relation extraction Jain et al. (2020). Document-level event extraction is important because the vast majority of real-world news and narrative texts describe multiple complex interacting events. Li et al. (2021), for example, find that the arguments of close to 40% of events in their Wikipedia-based dataset appear in sentences other than the event trigger sentence. Unfortunately, the document-level task setting poses significant challenges (Das et al., 2022). Documents can contain zero events or discuss multiple events of the same or similar types. Events and event arguments are likely to be mentioned many times throughout the document, requiring coreference resolution . Arguments can be shared across events. Information about different events can be intermingled.
@@
Note that event trigger-span identification is not required as part of the output; indeed, most datasets for the task do not include trigger-span annotations Soboroff (2023); Jain et al. (2020); Pavlick et al. (2016). On one hand, this makes sense: there is not generally a single trigger associated with each event, and the existence of intermingled event descriptions means that an event trigger may be associated with more than one event. Also, it is known that event individuation — the problem of distinguishing distinct events — is difficult even for people Gantt et al. (2023). On the other hand, event triggers, which are by definition lexical indicators of an event, should help with event type identification. In cases where they denote a (verb or noun) predicate for the event, they will serve as an anchor to localize argument extraction Ebner et al. (2020). Finally, consecutive triggers denoting different event types will be cues to segment the text into event-specific chunks.
@@
In this paper, we provide the first investigation of the usefulness of triggers for document-level event extraction. As described in Section 3, we treat Triggers as "rationales" that indicate the presence of an event and guide the system in finding the argument of the event. We analyze the usefulness of trigger rationales in multiple end-to-end and pipelined neural event extraction models for three document-level event extraction datasets. In particular, we measure performance when using a range of event triggers of varying quality (human-annotated, LLM-generated, keyword-based, and random) to perform event extraction. Additionally, we compare the results to in-context learning prompting baselines. We make available publicly all trigger annotations employed in our experiments [^2] GitHub Link (not available during anonymous review) including our manually annotated trigger-spans for the classic MUC dataset.
@@
We find that the usefulness of triggers depends, in part, on characteristics of the extraction task and its data, i.e. the density of relevant events in each document and the degree to which event arguments are localized with respect to the trigger. We also find that lower-quality, automatically identified triggers can be used as an alternative to human annotations. This is good news given the expense of obtaining manual annotations. In addition, we show that performance robustness in the face of degraded trigger quality can be achieved by making descriptive event information available to the extraction model. Perhaps surprisingly, we also find that the mere availability of triggers, even random ones, is important for prompt-based LLM approaches to the task. See Section 4.3 for access to our public data and code.
@@
Figure 2: Approach Overview: We examine four state-of-the art event extraction systems (e.g. TANL, Degree, GTT, GenIE). We run these models in two architectures, when possible, — pipeline and end-to-end (E2E). We also include comparisons to GPT-4o and GPT-4o-mini prompted via in-context few-shot learning (not shown). Models differ in the task-specific information document only (TANL, GTT), document+instructions (TANL +GenIE), document+instructions+template (Degree).
@@ 2 Related Work
Document-level event-based template filling/extraction was introduced in the Message Understanding Conferences (MUC) Sundheim (1991) as a task to identify occurrences of predefined event types and fill the template roles by extracting entity mentions from the text or categorizing aspects of the event. Annotation efforts to create such document-level datasets are limited in number and size due to their complex annotation instructions Sundheim (1991); Li et al. (2021); Soboroff (2023); Pavlick et al. (2016). Others have tried to collect such annotations automatically to create much larger datasets Jain et al. (2020); Tong et al. (2022), but the preference remains to benchmark against human-annotated text. Recent neural network systems tackle document-level template extraction using fine-tuning Hsu et al. (2022); Du et al. (2021), transfer learning from relevant tasks Wadden et al. (2019); Paolini et al. (2021), imitation learning Chen et al. (2023), and GPT-$n$ prompting Gao et al. (2023); Sun et al. (2024).
@@
Argument extraction is a subtask of event extraction that requires identifying fillers for the roles associated with the event given an event-type and its trigger-span Li et al. (2021); Ren et al. (2023); Ma et al. (2022). At the document level, various methods have been proposed. Retrieval-augmented methods Ren et al. (2023); Ma et al. (2022) retrieve relevant text from a training dataset before the extraction process, reducing the demand for model weights to memorize all forms and patterns useful for extraction. Conditional or constrained generation Li et al. (2021) forces the models to generate input document tokens corresponding to event argument mentions when filling natural sentences that verbalize a single template. To improve the global context especially when predicting multiple templates, Du et al. (2022) proposed to retrieve the previous template predictions dynamically to inform later prediction of the same documents. While a trigger and a linearized template are both used during an extraction, it is unclear if both are needed.
@@
3 Approach
Many NLP systems predict rationales in conjunction with outputs as a means of explaining them. In this study, we effectively treat triggers as rationales: a system is given, or predicts, event triggers, to then generate events and their arguments based on them. Previous studies have illustrated that NLP systems can leverage rationales in various ways. DeYoung et al. (2020) assert that rationales can function as explanations, helping end users evaluate a system’s trustworthiness. Yang et al. (2018b) show that providing rationales as extra supervision enhances a model’s capability to perform multi-hop reasoning. Most recently, Chen et al. (2022) demonstrate the potential of rationale-based methods to construct more resilient NLP systems. Given these benefits, we investigate whether triggers can offer similar advantages to document-level event extraction systems and prompting-based baselines.
Our approach for the evaluation is depicted in Figure 2. Specifically, we study triggers in conjunction with four recent approaches to document-level event extraction: TANL Paolini et al. (2021), GTT Du et al. (2021), Degree Hsu et al. (2022), and GenIE Li et al. (2021). In addition to adopting different underlying neural architectures, each espouses input requirements that differ in the amount of task-specific knowledge they expect. The models and input settings are described in Section 5. We consider both end-to-end and pipelined system architectures (Section 5.2) and provide comparisons to GPT-4o and GPT-4o-mini baselines prompted via in-context learning. Details on the experimental setup can be found in Section 6. Results are presented and analyzed in Section 7.
@@ 4 Datasets and Trigger Creation
Similar to previous work, we benchmark our results against the MUC Sundheim (1991), WikiEvents Li et al. (2021) and CMNEE Zhu et al. (2024) datasets. These are described next followed by the event trigger generation process.
4.1 Datasets Overview
@@
is a dataset introduced to study event extraction in English news articles reporting Latin American terrorist incidents. The dataset consists of document-template(s) pairs covering six event types. A document can contain no events or multiple events; similarly, an argument role can have no argument entities or multiple entities. Following previous work (Du et al., 2021; Wang et al., 2023a), we reduced the roles to be detected from 24 to 6 — event-type, perpetrator, perpetrator-organization, target, victim, weapon. MUC does not provide any trigger-span annotations, but we obtain Human (manually identified) trigger annotations as part of this work. The annotation process is described in Section 4.3.
@@
is a collection of English Wikipedia texts scraped from its “event" pages. The dataset consists of document-template(s) pairs covering 33 event types. In contrast to MUC, each role can have at most one entity. WikiEvents provides human (manually identified) trigger-span annotations.
@@
is a Chinese dataset of reports of military activity. The dataset consists of document-template(s) pairs covering eight event types; argument roles always have at least one entity. Since the dataset is in Chinese, we translated the input documents to English using GoogleTranslationAPI. We used HTML tags to delimit entity mentions before passing it through the translator to recover their translated correspondences. Because tags are sometimes dropped during translation, some mentions are unrecoverable. We filter out documents with such issues resulting in retaining about  60% of all documents across all splits. CMNEE provides human trigger-span annotations.
@@ 4.2 Dataset Analysis
The characteristics of the datasets vary considerably. Table 1 shows that WikiEvents has many events per document compared to MUC and CMNEE. MUC has few events on average per document whereas all of the WikiEvents documents and 62% of the CMNEE documents contain at least one event.
Finally, the localization of the argument mentions for event descriptions in the documents of each dataset varies tremendously — for WikiEvents, the arguments for each event are 1-2 tokens apart on average; for CMNEE, arguments for each event are generally within 8-9 tokens of each other; arguments for MUC events, on the other hand, are very spread out, spanning about 40 tokens.
@@ 4.3 Trigger Creation
We investigated four sources of trigger types: human, LLM, keyword, and random and provide details on how each is generated below. We will make publicly available all trigger sets employed in our experiments. [^3] GitHub Link (not available during anonymous review)
@@
human trigger spans are obtained by hiring annotators to identify the most informative trigger for each event in the document. They are the most expensive trigger type to obtain. WikiEvents and CMNEE provide one human-annotated trigger per event; MUC does not. As a result, we hired 9 annotators to identify all plausible trigger spans for each document-event pair in MUC. 57% of the document-template pairs in the dataset were annotated by at least two annotators. Average $\alpha$-Krippendorff and $F$-Kappa agreement scores across annotators of $0.530$ and $0.453$, respectively, indicate a moderate level of agreement. The annotation guidelines are provided in Appendix A.
@@
Keyword triggers are obtained by identifying a set of event-specific keywords at the corpus level and applying them heuristically for each document. For each event type $e\in E$, we identified documents that contained at least one event of type e to form a list $L_{\textit{e}}$ of words associated with e. We filtered each $L_{\textit{e}}$, keeping only the words that are plausible triggers for e, thus forming a list of keywords for each event type. We then heuristically picked entries $L_{\textit{e}}$ to annotate all events of type e. For example, “kill” frequents as a keyword for Attack events in MUC.
@@
We collected Random triggers by randomly selecting a span of 1-3 words from the document. They are obviously of poor quality and are used as a control to identify if the mere presence of triggers in the system input is helpful.
Examples of each trigger source are provided in Figure 1.
@@
GPT-4o-miniprompting .356 .514 .272 .256 .173 .487 .563 .671 .486
GPT-4oprompting .344 .562 .248 .266 .174 .570 .632 .565 .719
$\textsc{TANL}_{e2e}$514 .483 .550 .269 .185 .492 .400 .393 .408
$\textsc{DEGREE}_{e2e}$550 .610 .500 .241 .166 .433 .718 .681 .759
$\textsc{GTT}_{e2e}$388 .288 .596 .318 .396 .265
@@
GPT-4o-miniprompting .433 .604 .337 .355 .253 .594 .562 .671 .483
$\textsc{TANL}_{pipeline}$415 .364 .484 .433 .405 .464 .500 .548 .459
$\textsc{Degree}_{pipeline}$509 .531 .488 .274 .205 .410 .637 .621 .654
TANL+$\textsc{GENIE}_{pipeline}$464 .443 .486 .382 .336 .443 .599 .798 .479
$\textsc{TANL}_{e2e}$483 .515 .455 .284 .191 .548 .402 .428 .378
$\textsc{DEGREE}_{e2e}$563 .579 .548 .222 .144 .484 .680 .629 .738
$\textsc{GTT}_{e2e}$365 .268 .574 .326 .392 .279
@@
GPT-4o-miniprompting .413 .603 .314 .362 .250 .653 .578 .687 .499
$\textsc{TANL}_{pipeline}$290 .229 .394 .157 .100 .358 .420 .389 .455
$\textsc{Degree}_{pipeline}$541 .590 .499 .265 .200 .393 .622 .603 .643
TANL+$\textsc{GENIE}_{pipeline}$490 .439 .501 .388 .365 .413 .558 .555 .562
$\textsc{TANL}_{e2e}$435 .384 .500 .225 .141 .560 .387 .347 .437
$\textsc{DEGREE}_{e2e}$523 .545 .504 .271 .184 .514 .580 .517 .661
$\textsc{GTT}_{e2e}$354 .248 .617 .332 .373 .299
@@
GPT-4o-miniprompting .425 .603 .328 .353 .242 .647 .580 .689 .501
$\textsc{TANL}_{pipeline}$446 .490 .409 .277 .241 .325 .422 .438 .408
$\textsc{Degree}_{pipeline}$532 .623 .464 .241 .185 .345 .532 .647 .452
TANL+$\textsc{GENIE}_{pipeline}$470 .452 .489 .316 .268 .385 .628 .671 .590
$\textsc{TANL}_{e2e}$470 .446 .497 .232 .148 .534 .391 .365 .420
$\textsc{DEGREE}_{e2e}$523 .536 .511 .273 .183 .542 .601 .566 .642
$\textsc{GTT}_{e2e}$402 .308 .578 .324 .367 .291
@@
GPT-4o-miniprompting .408 .602 .309 .361 .259 .598 .528 .694 .426
$\textsc{TANL}_{pipeline}$109 .067 .312 .029 .007 .353 .175 .108 .455
$\textsc{Degree}_{pipeline}$538 .542 .535 .255 .192 .379 .589 .613 .566
TANL+$\textsc{GENIE}_{pipeline}$322 .264 .414 .120 .069 .449 .487 .392 .643
$\textsc{TANL}_{e2e}$402 .345 .482 .269 .183 .506 .340 .264 .478
$\textsc{DEGREE}_{e2e}$514 .492 .539 .270 .185 .501 .595 .561 .633
$\textsc{GTT}_{e2e}$304 .199 .651 .310 .311 .310
Table 2: Event extraction experiments. The F1-micro, recall and precision scores for the prompting baselines, end-to-end and pipeline systems on MUC, WikiEvents and CMNEE. Bolded scores refer to the best model across a dataset.. Scores in italics indicate the best model for a given a trigger setting.

### v2

@@ Abstract
Most existing work on event extraction has focused on sentence-level texts and presumes the identification of a trigger-span — a word or phrase in the input that evokes the occurrence of an event of interest. Event arguments are then extracted with respect to the trigger. Indeed, triggers are treated as integral to, and trigger detection as an essential component of, event extraction. In this paper, we provide the first investigation of the role of triggers for the more difficult and much less studied task of document-level event extraction. We analyze their usefulness in multiple end-to-end and pipelined transformer-based event extraction models for three document-level event extraction datasets, measuring performance using triggers of varying quality (human-annotated, LLM-generated, keyword-based, and random). We find that whether or not systems benefit from explicitly extracting triggers depends both on dataset characteristics (i.e. the typical number of events per document) and task-specific information available during extraction (i.e. natural language event schemas). Perhaps surprisingly, we also observe that the mere existence of triggers in the input, even random ones, is important for prompt-based in-context learning approaches to the task.
@@ 1 Introduction
Figure 1: Document-level event extraction example from MUC. Two events should be extracted from the given text, an Attack (doc on left) and a Kidnapping (doc on right). Note that only one mention (among many coreferent ones) per event argument is shown. Some event arguments are empty and some might have multiple fillers. The (duplicated) document depicts event triggers obtained from different annotation sources (labeled with pink background and the annotation ${}^{\text{and the annotation source}}$); more details in Section 3.
@@
In recent years, however, there has been increasing interest in document-level event extraction: given a document $d$, a set of event types $E$, and their associated arguments $A_{e\in E}$, produce one filled event template for each relevant event described in $d$ with event arguments recognized at the entity level, i.e., via a single representative mention. [^1] Other names for the task include template filling Jurafsky and Martin (2024), generalized template extraction (Chen et al., 2023), N-ary relation extraction Jain et al. (2020). Document-level event extraction is important because the vast majority of real-world news and narrative texts describe multiple complex interacting events. Li et al. (2021), for example, find that the arguments of close to 40% of events in their Wikipedia-based dataset appear in sentences other than the event trigger sentence. Unfortunately, the document-level task setting poses significant challenges (Das et al., 2022). Documents can contain zero events or discuss multiple events of the same or similar types. Events and event arguments are likely to be mentioned many times throughout the document, requiring coreference resolution. Arguments can be shared across events. Information about different events can be intermingled.
@@
Note that event trigger-span identification is not required as part of the output; indeed, most datasets for the task do not include trigger-span annotations Soboroff (2023); Jain et al. (2020); Pavlick et al. (2016). On one hand, this makes sense: a single event may have multiple triggers (or possibly none), and conversely a particular trigger text span may be associated with more than one event. Also, it is known that event individuation — the problem of distinguishing distinct events — is difficult even for human experts Gantt et al. (2023). On the other hand, event triggers, which are by definition lexical indicators of an event, should help with event type identification and systems may benefit from extracting them as an intermediate step. In cases where they denote a (verb or noun) predicate for the event, they will serve as an anchor to localize argument extraction Ebner et al. (2020). Finally, consecutive triggers denoting different event types will be cues to segment the text into event-specific chunks.
@@
In this paper, to the best of our knowledge, we provide the first investigation of the usefulness of triggers for document-level event extraction. As described in Section 3, we treat Triggers as “rationales" that indicate the presence of an event and guide the system in finding the argument of the event. We analyze whether using triggers improves event extraction performance in multiple end-to-end and pipelined transformer-based Vaswani et al. (2017) event extraction models for three document-level event extraction datasets. In particular, we measure event extraction performance for a range of event triggers of varying quality (human-annotated, LLM-generated, keyword-based, and random). Additionally, we compare the results to in-context learning prompting baselines. We publicly release all trigger annotations used in our experiments, [^2] GitHub Link including our manually annotated trigger-spans for the classic MUC dataset.
@@
We find that the usefulness of triggers depends, in part, on characteristics of the extraction task and its data, i.e., the density of relevant events in each document and the degree to which event arguments are localized with respect to the trigger. We also find that lower-quality, automatically identified triggers can be used as an alternative to human annotations. This is good news given the cost of obtaining manual annotations. In addition, we show that performance robustness in the face of degraded trigger quality can be achieved by making descriptive event information available to the extraction model. Perhaps surprisingly, we also find that the mere availability of triggers, even random ones, is important for prompt-based in-context learning approaches to the task. See Section 4.3 for access to our public data and code.
@@
Figure 2: Approach Overview: We examine four state-of-the art event extraction systems (e.g., TANL, Degree, GTT, GenIE). We run these models in two architectures, when possible, pipeline and end-to-end (E2E). We also include comparisons to GPT-4o and GPT-4o-mini prompted via in-context few-shot learning (not shown). Models differ in the task-specific information they provide—document only (TANL, GTT), document+instructions (TANL +GenIE), document+instructions+template (Degree).
@@ 2 Related Work
Document-level event extraction was introduced in the Message Understanding Conferences (MUC) Sundheim (1991) as a task to identify occurrences of predefined event types and fill the associated template roles by extracting entity mentions from the text or categorizing aspects of the event. Other event extraction datasets include RAMS Ebner et al. (2020), CMNEE Zhu et al. (2024), WikiEvents Li et al. (2021), and MAVEN Wang et al. (2020). We focus our analysis on MUC, WikiEvents, and CMNEE due to their long-document, multi-event nature and because their limited selection of event types enables clearer analysis of performance differences between our benchmarked systems. Document-level event extraction systems utilize drastically different techniques: earlier methods relied on pattern matching Riloff (1993); Riloff and Jones (1999), while more modern approaches use deep learning to represent event components such as triggers and entities Wadden et al. (2019); Wang et al. (2023c); Xu et al. (2021). Most recently, the task has been framed as sequence-to-sequence generation (i.e. TANL Paolini et al. (2021), GTT Du et al. (2021) and Degree Hsu et al. (2022)) built upon powerful transformer Vaswani et al. (2017) language models. Our work focuses primarily on such methods due to the flexibility of sequence-to-sequence frameworks Paolini et al. (2021) and because we believe future methods will increasingly utilize constantly improving LLMs.
@@
Argument extraction is a subtask of event extraction that requires identifying fillers for the roles associated with the event given an event type as well as its trigger span Li et al. (2021); Ren et al. (2023); Ma et al. (2022). Argument extractors are commonly employed in event extraction systems that segment the task into trigger/event detection and typing followed by argument extraction. We primarily focus on TANL Paolini et al. (2021) and GenIE Li et al. (2021) that model the task via sequence generation, though other methods that rely on careful manipulation and learning of embeddings also exist Chen et al. (2015); Yang et al. (2019).
3 Approach
Many NLP systems predict rationales in conjunction with outputs as a means of explaining them. In this study, we effectively treat triggers as rationales: a system is given, or predicts, event triggers that may guide identification of associated arguments. Previous studies have illustrated that NLP systems can leverage rationales in various ways. DeYoung et al. (2020) assert that rationales can function as explanations, helping end users evaluate a system’s trustworthiness. Yang et al. (2018b) show that providing rationales as extra supervision enhances a model’s capability to perform multi-hop reasoning. Most recently, Chen et al. (2022) demonstrate the potential of rationale-based NLP systems on adversarial inputs. Given these benefits, we investigate whether triggers can offer similar advantages to document-level event extraction systems and prompting-based baselines.
Our approach for the evaluation is depicted in Figure 2. Specifically, we study triggers in conjunction with four recent sequence-to-sequence transformer-based approaches to document-level event extraction: TANL Paolini et al. (2021), GTT Du et al. (2021), Degree Hsu et al. (2022), and GenIE Li et al. (2021). In addition to adopting different underlying neural architectures, each has input requirements that differ in the amount of task-specific knowledge they expect. The models and input settings are described in Section 5. We consider both end-to-end and pipelined system architectures (Section 5.2) and provide comparisons to GPT-4o and GPT-4o-mini baselines prompted via in-context learning. Details on the experimental setup can be found in Section 6. Results are presented and analyzed in Section 7.
We specifically investigate the following:
  1. 1.
    Do corpora characteristics, such as the typical number of events and the distribution of event arguments across a document, influence the utility of triggers? For example, triggers might be useful anchors for argument extraction when there are many events with arguments located in close proximity to triggers.
  2. 2.
    Can we view triggers as grounding information for the event extraction task, and if so, could other task/corpus specific information be utilized in place of triggers such that trigger-free methods can match trigger-based methods? For example, can methods that ingest fine-grained event schema details without explicitly identifying triggers during extraction be as performant as methods explicitly learning trigger identification? If grounding information beyond triggers is effective and can be easily constructed, we may be able to bypass expensive human trigger annotation efforts during event extraction dataset creation.
  3. 3.
    To further minimize manual annotation, perhaps only providing in-context examples to a modern LLM can achieve reasonable performance. Alternatively, the many subtasks subsumed in event extraction (event detection, entity detection, entity and event typing) may still be too complex for LLMs. We also explore how triggers impact in-context learning methods.
  4. 4.
    Event extraction methods are generally either pipelined (extract a trigger, then extract a single event’s arguments conditioned on a trigger) or end-to-end (the model learns trigger and argument extraction simultaneously). Which performs better? Does the source of the trigger annotations impact performance between the two paradigms? While our benchmark methods by default fit exclusively in one category, slight modifications allow single systems to model both paradigms such that we can center our analyses around the paradigm used.
@@ 4 Datasets and Trigger Creation
Similar to previous work, we perform our analyses on the MUC Sundheim (1991), WikiEvents Li et al. (2021) and CMNEE Zhu et al. (2024) datasets. These are described next, followed by the event trigger generation process.
4.1 Datasets
@@
is a dataset introduced to study event extraction in English news articles reporting Latin American terrorist incidents. The dataset consists of document-template(s) pairs covering six event types. A document can contain no events or multiple events; similarly, an argument role can have no argument entities or multiple entities. Following previous work (Du et al., 2021; Wang et al., 2023a), we restricted event templates to contain 6 out of the original 24 roles — event-type, perpetrator, perpetrator-organization, target, victim, weapon. MUC does not provide any trigger-span annotations, but we obtain Human (manually identified) trigger annotations as part of this work. The annotation process is described in Section 4.3.
@@
is a collection of English Wikipedia texts scraped from its “event" pages. The dataset consists of document-template(s) pairs covering 33 event types. In contrast to MUC, each role is filled by at most one entity. WikiEvents provides human (manually identified) trigger-span annotations.
@@
is a Chinese dataset of reports of military activity. The dataset consists of document-template(s) pairs covering eight event types; argument roles always have at least one entity filler. Since the dataset is in Chinese, we translated the documents to English using the Google Translate API. We used HTML tags to delimit entity mentions before passing it through the translator to recover their translated correspondences. Because tags are sometimes dropped during translation, some mentions are unrecoverable. We filter out documents with such issues resulting in retaining about  60% of all documents. CMNEE provides human trigger-span annotations.
@@ 4.2 Dataset Analysis
The characteristics of the datasets vary considerably. Table 1 shows that WikiEvents has many events per document compared to MUC and CMNEE. MUC has few events on average per document (many documents have no associated events) whereas all of the WikiEvents documents and 62% of the CMNEE documents contain at least one event.
Finally, the localization of the argument mentions for event descriptions in the documents of each dataset varies tremendously — for WikiEvents, the arguments for each event are 1-2 tokens apart [^3] We used nltk.word_tokenize to tokenize documents. For each event, we identify the index of the each of its arguments’ first token, then compute the standard deviation of these indices. on average; for CMNEE, arguments for each event are generally within 8-9 tokens of each other; arguments for MUC events, on the other hand, are very spread out, spanning about 40 tokens.
Ideally the documents in our datasets would both be long and have event arguments widely dispersed throughout the texts, but such a dataset does not exist at the time of writing. Our selection of MUC, WikiEvents, and CMNEE as a group nevertheless still allows us to perform meaningful analyses of these characteristics.
@@ 4.3 Trigger Creation
We investigated four sources of trigger annotation: human, LLM, keyword, and random and provide details on how each is generated below. Examples of each trigger source are provided in Figure 1. We will make publicly available all trigger sets employed in our experiments. [^4] GitHub Link (not available during anonymous review)
@@
human trigger spans are obtained by hiring annotators to identify the most informative trigger for each event in the document. They are the most expensive trigger type to obtain. WikiEvents and CMNEE provide one human-annotated trigger per event; MUC does not. As a result, we hired 9 annotators to identify all plausible trigger spans for each document-event pair. 57% of the document-template pairs in the dataset were annotated by at least two annotators. Average $\alpha$-Krippendorff and $F$-Kappa agreement scores across annotators of $0.530$ and $0.453$, respectively, indicate a moderate level of agreement. The annotation guidelines are provided in Appendix A.
@@
Keyword triggers are obtained by identifying a set of event-specific keywords at the corpus level and applying them heuristically for each document. For each event type $e\in E$, we identified documents that contained at least one event of type e to form a list $L_{\textit{e}}$ of words associated with e. We manually filtered each $L_{\textit{e}}$ producing lists $\hat{L}_{\textit{e}}$ that contain only the words that are plausible triggers for e (i.e., “exploded" or “bombed" for a bombing event), thus forming a list of keywords for each event type. We then use the entries from $\hat{L}_{\textit{e}}$ to annotate all events of type e.
@@
We also collected Random triggers by randomly selecting a span of 1-3 words from the document. They are obviously of poor quality and are used as a control to identify whether the mere presence of triggers in the system input is helpful.
@@
GPT-4o-miniprompting .$434_{\pm.01}.$$325_{\pm.02}.$$565_{\pm.00}.$
GPT-4oprompting .$391_{\pm.01}.$$339_{\pm.02}.$$635_{\pm.00}.$
$\textsc{{TANL}}_{e2e}$$559_{\pm.01}.$$361_{\pm.02}.$$438_{\pm.00}.$
$\textsc{{DEGREE}}_{e2e}$$597_{\pm.00}.$$248_{\pm.01}.$$789_{\pm.00}.$
$\textsc{{GTT}}_{e2e}$$454_{\pm.01}.$$318_{\pm.00}.$
@@
GPT-4o-miniprompting .$438_{\pm.02}.$$364_{\pm.02}.$$564_{\pm.00}.$
GPT-4oprompting .$405_{\pm.01}.$$379_{\pm.02}.$$667_{\pm.00}.$
$\textsc{{TANL}}_{pipeline}$$460_{\pm.01}.$$471_{\pm.02}.$$596_{\pm.01}.$
$\textsc{{DEGREE}}_{pipeline}$$524_{\pm.01}.$$357_{\pm.01}.$$723_{\pm.00}.$
TANL+$\textsc{GENIE}_{pipeline}$$464_{\pm.01}.$$382_{\pm.01}.$$599_{\pm.00}.$
$\textsc{{TANL}}_{e2e}$$530_{\pm.01}.$$337_{\pm.02}.$$410_{\pm.00}.$
$\textsc{{DEGREE}}_{e2e}$$566_{\pm.01}.$$339_{\pm.02}.$$782_{\pm.00}.$
$\textsc{{GTT}}_{e2e}$$429_{\pm.01}.$$326_{\pm.00}.$
@@
GPT-4o-miniprompting .$416_{\pm.02}.$$368_{\pm.02}.$$580_{\pm.00}.$
GPT-4oprompting .$414_{\pm.01}.$$383_{\pm.02}.$$659_{\pm.00}.$
$\textsc{{TANL}}_{pipeline}$$305_{\pm.01}.$$207_{\pm.01}.$$543_{\pm.00}.$
$\textsc{{DEGREE}}_{pipeline}$$567_{\pm.01}.$$335_{\pm.01}.$$715_{\pm.00}.$
TANL+$\textsc{GENIE}_{pipeline}$$490_{\pm.01}.$$388_{\pm.01}.$$558_{\pm.00}.$
$\textsc{{TANL}}_{e2e}$$507_{\pm.00}.$$267_{\pm.02}.$$440_{\pm.00}.$
$\textsc{{DEGREE}}_{e2e}$$557_{\pm.00}.$$313_{\pm.02}.$$703_{\pm.00}.$
$\textsc{{GTT}}_{e2e}$$439_{\pm.02}.$$332_{\pm.00}.$
@@
GPT-4o-miniprompting .$429_{\pm.02}.$$361_{\pm.02}.$$582_{\pm.00}.$
GPT-4oprompting .$420_{\pm.01}.$$389_{\pm.02}.$$659_{\pm.00}.$
TANLpipeline .$525_{\pm.02}.$$306_{\pm.02}.$$640_{\pm.00}.$
DEGREEpipeline .$532_{\pm.01}.$$241_{\pm.01}.$$532_{\pm.00}.$
TANL+$\textsc{GENIE}_{pipeline}$$470_{\pm.01}.$$316_{\pm.01}.$$628_{\pm.00}.$
TANLe2e .$495_{\pm.00}.$$292_{\pm.02}.$$443_{\pm.00}.$
DEGREEe2e .$523_{\pm.01}.$$273_{\pm.01}.$$601_{\pm.00}.$
GTTe2e .$455_{\pm.01}.$$324_{\pm.00}.$
@@
GPT-4o-miniprompting .$412_{\pm.01}.$$369_{\pm.02}.$$533_{\pm.01}.$
GPT-4oprompting .$402_{\pm.00}.$$359_{\pm.02}.$$533_{\pm.01}.$
TANLpipeline .$116_{\pm.01}.$$059_{\pm.01}.$$223_{\pm.00}.$
DEGREEpipeline .$538_{\pm.01}.$$255_{\pm.01}.$$589_{\pm.00}.$
TANL+$\textsc{GENIE}_{pipeline}$$322_{\pm.01}.$$120_{\pm.01}.$$487_{\pm.00}.$
TANLe2e .$473_{\pm.01}.$$343_{\pm.02}.$$394_{\pm.00}.$
DEGREEe2e .$514_{\pm.01}.$$270_{\pm.02}.$$595_{\pm.01}.$
GTTe2e .$304_{\pm.01}.$$310_{\pm.00}.$
Table 2: Event extraction experiments. Average F1-micro scores on 5-fold cross validation for the prompting baselines, end-to-end and pipeline systems on MUC, WikiEvents and CMNEE. The colored cells indicate the best performing model, for a given dataset, with the darkest color indicating the highest F1 score.